# Quantum reservoir probing of quantum phase transitions

### v2

@@ RESULTS
Figure 3: Detection of the quantum phase transition in the ANNNI model using the QRP. a Spatiotemporal representation of the estimation performance when employing $\langle\sigma^{x}_{i}(t)\rangle$ in the nonintegrable ANNNI chain characterized by the interaction parameter $\kappa=0.5$. The system size is $N=49$, and the time evolution is calculated with a bond dimension of $\chi=128$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{0<t\leq 15\}$, where $\Lambda$ contains the central eleven sites. The upper band represents the schematic phase diagram, where the ferromagnetically ordered and quantum disordered phases are separated by the quantum critical point at approximately $g\simeq 1.6$.
To illustrate the applicability of our QRP framework to nonintegrable models, we extend the model to the anisotropic next-nearest-neighbor Ising (ANNNI) chain [41, 42, 43, 44]. The Hamiltonian is given by
@@
where we take $J=1$ and $\kappa$ represents the strength of the next-nerarest-neighbor interaction. The latter term makes the system nonintegrable by introducing the four-body interactions for the quasiparticles derived by the Jordan-Wigner transformation. This model hosts a rich phase diagram in the $\kappa$-$g$ plane [45]. Our investigation focuses on the case of $\kappa=0.5$, where the quantum phase transition occurs at $g_{c}\simeq 1.6$ with the Ising universality, separating ferromagnetically ordered ($g<g_{c}$) and quantum disordered ($g>g_{c}$) phases. The initial state is prepared in the same manner as Eq. (6).
@@
Figure 3a displays the spatiotemporal representation of $R^{2}_{i}(t)$ obtained from $\langle\sigma^{x}_{i}(t)\rangle$, revealing a scenario similar to the transverse-field Ising model. For the majority of sites and times, the local spin $\langle\sigma_{i}^{x}(t)\rangle$ mirrors the input value $s_{k}$ employed in the local quantum quench, as evidenced by the nonzero values of $R^{2}_{i}(t)$. Above the critical field at $g_{c}\simeq 1.6$, particularly for $g=2.2$ in the right panel, the effects of the local manipulation propagate primarily through the excitation wavefronts, while the local spin oscillations exhibit weaker dependence on the local quench with small $R^{2}_{i}(t)$. Conversely, below the critical point the opposite behavior is noted in the left panel for $g=1.0$. Furthermore, in the vicinity of the quantum critical point, the influence of the local quench manifest itself predominantly along the wavefronts, while in the other regions it is suppressed by the enhanced fluctuations, as demonstrated by the emergence of a discernible dark zone between the wavefronts. These fundamental dependencies of the local spin operators on the local quench retain a similarity to those observed in the integrable model presented in Fig. 2a.
@@
In Fig. 3b, we present the mean value $\overline{R^{2}}$, calculated over the indices $(i,t)$ corresponding to the central eleven sites within the time frame $0<t\leq 15$ (see the Supplemental Information for other subsets). Analogous to the observation in Fig. 2d, $\overline{R^{2}}$ demonstrates a pronounced decrease around the quantum critical point at $g\simeq 1.6$, precisely locating the boundary between the different spatiotemporal distribution patterns of $R^{2}_{i}(t)$ in the ferromagnetically ordered phase and the quantum disordered phase. This observation confirms that, even in nonintegrable systems, $R^{2}_{i}(t)$ is an effective marker for detecting quantum phase transitions.
Spin-1 XXZ model
@@
The preceding sections have demonstrated the effectiveness of our QRP in identifying quantum phase transitions, particularly between conventional phases where the internal physics is well-characterized and understood within the Landau-Ginzburg-Wilson theory. Building on this foundation, we now aim to broaden our exploration to more complex quantum phenomena. In particular, we apply the QRP framework to the detection of topological quantum phase transitions, which lack local order parameters that can distinguish adjacent phases. This raises the question of whether the QRP framework, which solely utilizes a single-site spin operator, can still be effective in identifying these transitions beyond the Landau-Gintzburg-Wilson paradigm.
@@
Here, we study the spin-$1$ XXZ model with open boundaries [46, 47], which is defined by
$\mathcal{H}=\sum_{i}J(S^{x}_{i}S^{x}_{i+1}+S^{y}_{i}S^{y}_{i+1})+\lambda S^{z}_{i}S^{z}_{i+1},$ (8)
@@
where $S^{\alpha}_{i}$ denotes the $\alpha$-component of spin-$1$ operator at site $i$ and $\lambda$ represents the Ising anisotropy; we set the interaction strength $J=1$. This model is renowned for hosting a prototypical symmetry-protected topological phase, known as the Haldane phase, observed for the anisotropy within $0<\lambda<\lambda_{c}$ with $\lambda_{c}\simeq 1.18$ [48, 49, 50]. The Haldane phase is distinguished by its protection under a $Z_{2}\times Z_{2}$ symmetry, and is characterized by nonlocal string order parameters [51, 52, 53]. With increasing the anisotropy, the topological quantum phase transition into the topologically trivial spin-$1$ Néel state occurs at the quantum critical point, which belongs to the two-dimensional (2D) Ising universality class [49, 54, 55].
The QRP framework can be naturally extended to the spin-$1$ XXZ model. Here, $|\Phi\rangle$ is initialized as the Néel state, and the input spin state $|\psi^{\prime}(s_{k})\rangle$, the spin-$1$ extension of Eq. (2), is prepared by applying a local magnetic field $\bm{h}^{\prime}\propto(s_{k}-1,0,-\sqrt{s_{k}(2-s_{k})})$. Explicitly, the resultant state $|\Psi^{\mathrm{in}}_{k}\rangle$ is expressed as:
$|\Psi^{\mathrm{in}}_{k}\rangle=\mid\uparrow\downarrow\uparrow\downarrow\dots\uparrow\downarrow\rangle\otimes|\psi^{\prime}(s_{k})\rangle\otimes\mid\downarrow\uparrow\downarrow\uparrow\dots\downarrow\uparrow\rangle.$ (9)
Following this setup, the impact of the local quench on $\langle S^{x}_{i}(t)\rangle$ is evaluated through the estimation performance $R^{2}_{i}(t)$, analogous to the previous cases.
@@
Figure 4: Probing the topological quantum phase transition in the spin-1 XXZ model using the QRP. a Spatiotemporal representation of the estimation performance $R_{i}^{2}(t)$ employing $\langle S^{x}_{i}(t)\rangle$ for a system with $N=49$ sites. The time evolution is calculated with a bond dimension of $\chi=128$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{10<t\leq 20\}$, where $\Lambda$ represents the entire system. The upper bar illustrates the phase diagram of $\mathcal{H}(\lambda)$ with schematics composed of virtual $S=1/2$ spins.
@@
In Fig. 4a, we present the estimation performance $R^{2}_{i}(t)$ under the Hamiltonian $\mathcal{H}$ for three different anisotropies: $\lambda=1.0$ in the Haldane phase, $\lambda=1.3$ in the Néel phase, and $\lambda=1.18$ at the vicinity of the quantum critical point. The spread of nonzero $R^{2}_{i}(t)$ values indicates that the operator $S_{i}^{x}(t)$ is reflective of the local quantum quench at various spatiotemporal points, similar to what is observed in $S=1/2$ spin systems. Following an initial transient period, $R^{2}_{i}(t)$ shows a disturbed ripple-like pattern in the Néel phase, reminiscent of those in the ordered phases in Figs. 2a and 2b. In contrast, in the Haldane phase, $R^{2}_{i}(t)$ becomes more uniform, presumably due to the nonlocal sharing of quenching effects via the underlying singlet pairs. Around the critical point, a notable trend is observed wherein $R^{2}_{i}(t)$ progressively declines over time. This precisely mirrors the trend observed at the quantum critical point in conventional quantum phase transitions (Figs. 2a and 3a).
@@
Figure 4b presents the mean value $\overline{R^{2}}$, which effectively distinguishes these three distinct distributions of $R^{2}_{i}(t)$. $\overline{R^{2}}$ is computed as the average of $R^{2}_{i}(t)$ for the entire system within the time interval $10<t\leq 20$. A key observation in Fig. 4b is the pronounced dip at approximately $\lambda\simeq\lambda_{c}$, signifying the suppression of the impact of local quench. This dip thus marks the quantum critical point characterized by enhanced fluctuations, clearly delineating the boundary between the topological Haldane phase and the trivial Néel phase. Consequently, despite the absence of local order parameters in the Haldane phase, our QRP, based on the local quench and local observables, demonstrates itself as a potent instrument for detecting the topological quantum phase transition.
Extended Su-Schrieffer-Heeger model
Finally, we analyze the extended Su-Schrieffer-Heeger (eSSH) model [56, 57, 58, 59, 60, 61]:
$\mathcal{H}=$ $J\sum_{i=1}^{N/2}(\sigma^{x}_{2i-1}\sigma^{x}_{2i}+\sigma^{y}_{2i-1}\sigma^{y}_{2i}+\delta\sigma^{z}_{2i-1}\sigma^{z}_{2i})$
$+$ $J^{\prime}\sum_{i=1}^{N/2-1}(\sigma^{x}_{2i}\sigma^{x}_{2i+1}+\sigma^{y}_{2i}\sigma^{y}_{2i+1}+\delta\sigma^{z}_{2i}\sigma^{z}_{2i+1}).$ (10)
Here, $J$ and $J^{\prime}$ represent the strength of alternating nearest-neighbor interactions, and $\delta$ denotes the anisotropy in the $z$ direction; we set $J=1$. Our analysis focuses on the scenario at $\delta=0.5$, where a topological phase transition occurs around $J^{\prime}_{c}\simeq 1$, typically examined through many-body topological invariants [62, 60, 61]. For $J^{\prime}<J^{\prime}_{c}$, the trivial dimer phase manifests, with spins forming singlet pairs with their nearest neighbors, while for $J^{\prime}>J^{\prime}_{c}$, the topological phase emerges, characterized by the emergence of edge states. Importantly, these phases are entirely topological and devoid of local orderings on both sides of the quantum critical point, unlike the spin-1 XXZ model, which exhibits the local Néel order on one side.
Figure 5: Detection of the topological quantum phase transition in the eSSH model using the QRP. a Color map of the estimation performance $R_{i}^{2}(t)$ employing $\langle\sigma^{x}_{i}(t)\rangle$ in the eSSH model with $\delta=0.5$. The system size is $N=48$, and the time evolution is calculated with a bond dimension of $\chi=128$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{8<t\leq 20\}$, where $\Lambda$ contains the central $24$ sites. The upper band represents the phase diagram, where the topological and trivial phases are separated by the quantum critical point at approximately $J^{\prime}\simeq 1.0$.
The initial state $|\Psi_{k}\rangle$ is prepared as the ground state of $\mathcal{H}$ in the topological phase with $J^{\prime}=6.0$, subject to a weak magnetic field $\bm{h}\propto(2s_{k}-1,0,-2\sqrt{s_{k}(1-s_{k})})$ applied locally to one of the central sites for encoding input information. To ensure that any variations in $|\Psi_{k}\rangle$ across different instances arise solely from differences in $s_{k}$, a magnetic field is applied at both edge spins along the $+z$ direction; this edge field resolves the ground state degeneracy associated with edge degrees of freedom. Both central and edge fields are set to a magnitude of $0.1$ and deactivated following state preparation. Although the prepared state is not a product state and thus not representable in the form of Eq. (1), the effects of the local field are essentially confined to the central region, thereby presenting a virtually identical scenario to the local quench operation. Commencing from this initial state, we monitor the dynamics of the local spin $\langle\sigma_{i}^{x}(t)\rangle$ across multiple cases with different $s_{k}$. The resultant discrepancies in the observed dynamics are attributed to the initial localized manipulation parameterized by $s_{k}$, analogous to prior models.
@@
Figure 5a illustrates the estimation performance $R^{2}_{i}(t)$ under the Hamiltonian $\mathcal{H}$ for three different values of $J^{\prime}$. In all scenarios, $R^{2}_{i}(t=0)$ acquires nonzero values only around the central region, confirming that the information regarding $s_{k}$ is indeed introduced locally through the application of the local magnetic field. During the subsequent time evolution, the effects of the initial localized manipulation begin to propagate throughout the system, where the distinctions between the phases become evident. Notably, in proximity to the quantum critical point at $J^{\prime}=1.0$, $R^{2}_{i}(t)$ is markedly suppressed, displaying a discernibly dark spatiotemporal region in Fig. 5a. This behavior contrasts with the topological phase at $J^{\prime}=1.3$, which exhibits a ripple-like pattern, and the trivial phase at $J^{\prime}=0.7$, characterized by a strip-shaped pattern. In Fig. 5b, we illustrate the $J^{\prime}$ dependence of $\overline{R^{2}}$, defined by averaging $R^{2}_{i}(t)$ over the central $24$ sites within the time frame $8<t\leq 20$. Consistent with previous models, $\overline{R^{2}}$ exhibits a pronounced dip at the quantum critical point, which precisely delineates the boundary between the trivial and topological phases. Indeed, although the transition is entirely topological and the local orderings are absent, the quantum fluctuations are consistently enhanced at the critical point. The QRP locally detect such fluctuations as a witness of the topological quantum phase transition.
@@ DISCUSSION
To summarize, we have proposed the QRP as an versatile tool for identifying the quantum phase transitions in a variety of quantum systems, including integrable, nonintegrable, and topological systems. The QRP selectively evaluates the effects of the local quench in a manner similar to the pump-probe paradigm. Here, the parametrized local quantum quench, acting as the “pump”, injects random disturbance into the dynamics. The estimation performance of this random input parameter, through the linear transformation of the single-site operator, serves as the “probe”. This pump-probe-like process informationally discerns the underlying relationship between the local quench and the resultant dynamics of operators. Using the QRP, we have observed that the local quench operation exerts varied influence depending on the nature of the quantum phases. At the quantum critical points, the introduced disturbance becomes subordinate to the maximally enhanced quantum fluctuations. A noticeable decline in the estimation performance observed therein serves as a marker for the transition between distinct quantum phases. These findings have established that local, nonequilibrium excitations bear the signatures of quantum phase transitions concerning global, equilibrium many-body ground states. The crux of our methodology lies in the suppression of local quenching effects on local dynamics due to amplified quantum fluctuations, the rigorous justification of which is left for future study.
@@
Our investigations have focused on 1D quantum spin systems; however, the QRP methodology is applicable to a broader range of quantum systems in more than one dimension as well. Indeed, the conformal field theories governing quantum phase transitions in two and higher-dimensional systems exhibit significant differences from their 1D counterparts [63]. These differences manifest themselves as changes in nature of excitations, intrinsic dynamics, and critical behaviors near quantum critical points, all of which would be detectable through the QRP. Moreover, geometric frustrations in high-dimensional quantum spin systems, coupled with the inherent quantum fluctuations, give rise to a plethora of topological quantum phase transitions in a variety of systems. For instance, in certain systems, such transitions are associated with the emergence of anyons, which have promising potential for applications in topological quantum computing [64]. The examination of how anyons respond to local quantum operations through the QRP holds significant importance in light of the aforementioned applications. We leave the high-dimensional applications of the QRP for future investigations.
@@
Finally, let us remark on the design flexibility of the QRP. A fundamental advantage of employing local quantum quenches, in contrast to global ones, lies in their capability to access and elucidate quantum phenomena induced by local operations, such as specific classes of excitations. Indeed, the concept of excitation in quantum systems is intrinsically linked to the relative relationship between local manipulations and initial states; thus, identical local quench operations can lead to varying outcomes based on the differences in the initial states. When one possesses prior knowledge regarding the physics of the targeted system, especially for discerning partially understood phenomena, the optimal selections for the initial state, local quantum quench, and read-out operators in the QRP become intuitively evident. In contrast, the true value of the QRP framework becomes markedly pronounced in the exploration of completely unknown quantum systems. Therein, the underlying physics remains undisclosed, rendering the interpretation of the complex dynamics observed a formidable task. Our QRP introduces a systematic methodology to address this complexity, categorizing operators by meticulously investigating their dependence on various local operations through a scan of the aforementioned conditions of the QRP. Furthermore, the experimental simplicity of the QRP is noteworthy, as it necessitates solely single-site measurements during the initial stages of the dynamics. Consequently, we believe that the QRP presents a promising methodology for eliciting novel insights into a broad spectrum of exotic quantum many-body phenomena, extending beyond quantum phase transitions.

### v3

@@ RESULTS
To illustrate the applicability of our QRP framework to nonintegrable systems, we extend the model to the anisotropic next-nearest-neighbor Ising (ANNNI) chain [41, 42, 43, 44]. The Hamiltonian is given by
@@
where $J$ and $\kappa$ represent the strengths of the nerarest-neighbor and the next-nerarest-neighbor interactions, respectively; we take $J=1$. The latter term makes the system nonintegrable by introducing the four-body interactions for the quasiparticles derived by the Jordan-Wigner transformation. This model hosts a rich phase diagram in the $\kappa$-$g$ plane [45]. Our investigation focuses on the case of $\kappa=0.5$, where the quantum phase transition occurs at $g=g_{c}$ with the Ising universality, separating ferromagnetically ordered ($g<g_{c}$) and quantum disordered ($g>g_{c}$) phases. In our finite-size system with open boundary conditions, we determined $g_{c}\simeq 1.7$ utilizing the density matrix renormalization group (DMRG) method. The initial state is prepared in the same manner as Eq. (6).
@@
Figure 3a displays the spatiotemporal representation of $R^{2}_{i}(t)$ obtained from $\langle\sigma^{x}_{i}(t)\rangle$, revealing a scenario similar to the transverse-field Ising model. For the majority of sites and times, the local spin $\langle\sigma_{i}^{x}(t)\rangle$ mirrors the input value $s_{k}$ employed in the local quantum quench, as evidenced by the nonzero values of $R^{2}_{i}(t)$. Above the critical field at $g_{c}\simeq 1.7$, particularly for $g=2.6$ in the right panel, the effects of the local manipulation propagate primarily through the excitation wavefronts, while the local spin oscillations exhibit weaker dependence on the local quench with small $R^{2}_{i}(t)$. Conversely, below the critical point the opposite behavior is noted in the left panel for $g=1.0$. Furthermore, in the vicinity of the quantum critical point, the influence of the local quench manifests itself predominantly along the wavefronts, while in the other regions it is suppressed by the enhanced quantum fluctuations, as demonstrated by the emergence of a discernible dark zone between the wavefronts. These fundamental dependencies of the local spin operators on the local quench retain a similarity to those observed in the integrable model presented in Fig. 2a.
@@
In Fig. 3b, we present the mean value $\overline{R^{2}}$, calculated over the indices $(i,t)$ corresponding to the central seven sites within the time frame $0<t\leq 5$ (see the Supplementary Information for other subsets). Analogous to the observation in Fig. 2d, $\overline{R^{2}}$ demonstrates a pronounced decrease around the quantum critical point at $g_{c}\simeq 1.7$, precisely locating the boundary between the different spatiotemporal distribution patterns of $R^{2}_{i}(t)$ in the ferromagnetically ordered phase and the quantum disordered phase. This observation confirms that, even in nonintegrable quantum systems, $R^{2}_{i}(t)$ functions as an effective marker for detecting quantum phase transitions.
Cluster model
@@
The preceding sections have demonstrated the effectiveness of our QRP in identifying quantum phase transitions, particularly between conventional phases where the internal physics is well-characterized and understood within the Landau-Ginzburg-Wilson theory. Building on this foundation, we now aim to broaden our exploration to more complex quantum phenomena. In particular, we apply the QRP framework to the detection of topological quantum phase transitions, which lack local order parameters that can distinguish adjacent phases. This raises the question of whether the QRP framework, which solely utilizes a single-site spin operator, can still be effective in identifying these transitions beyond the Landau-Ginzburg-Wilson paradigm.
@@
Here, we study the cluster model with open boundaries [46, 47, 48], which is defined by
$\mathcal{H}=-\sum_{i}J_{zz}\sigma_{i}^{z}\sigma_{i+1}^{z}+\sum_{i}J_{zxz}\sigma_{i}^{z}\sigma_{i+1}^{x}\sigma_{i+2}^{z},$ (8)
@@
where $J_{zz}$ and $J_{zxz}$ denote the Ising and cluster interaction strengths, respectively; we set $J_{zz}=1$. This model is renowned for hosting a symmetry-protected topological (SPT) phase, known as the cluster state, which emerges for $J_{zz}<J_{zxz}$. The SPT phase is distinguished by its protection under a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{T}$ symmetry, and is characterized by nonlocal string order parameters [49, 50, 51, 52, 53]. As the strength of the cluster interaction diminishes, the quantum phase transition occurs from the SPT phase to the topologically trivial, ferromagnetically ordered phase at the quantum critical point $J_{zxz}=1$. We aim to detect this transition through the dynamical signature in the QRP, by preparing the initial state in Eq. (6) and monitoring the dynamics of the local spin $\sigma_{i}^{x}(t)$.
@@
Figure 4: Probing the quantum phase transition between the ferromagnetically ordered and SPT phases in the cluster model using the QRP. a Spatiotemporal representation of the estimation performance $R_{i}^{2}(t)$ employing $\langle\sigma^{x}_{i}(t)\rangle$ for a system with $N=49$ sites. The time evolution is calculated with a bond dimension of $\chi=256$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{0<t\leq 6.5\}$, where $\Lambda$ represents the central eleven spins. The upper bar illustrates the phase diagram.
@@
In Fig. 4a, we present the estimation performance $R^{2}_{i}(t)$ under the Hamiltonian $\mathcal{H}$ for three different strengths of the cluster interaction: $J_{zxz}=0.6$ in the ferromagnetically ordered phase, $J_{zxz}=1.4$ in the SPT phase, and $J_{zxz}=1.0$ at the quantum critical point. Consistent with the previous topologically trivial models, the spread of nonzero $R^{2}_{i}(t)$ values in the SPT phase indicates that the operator $\sigma_{i}^{x}(t)$ is reflective of the local quantum quench at various spatiotemporal points. Importantly, at the quantum critical point, a notable trend is observed wherein $R^{2}_{i}(t)$ is suppressed following the passage of the wavefronts. This precisely mirrors the trend observed at the quantum critical point in conventional quantum phase transitions (Figs. 2d and 3b).
@@
Figure 4b presents the mean value $\overline{R^{2}}$, which effectively distinguishes these three distinct distributions of $R^{2}_{i}(t)$. $\overline{R^{2}}$ is computed as the average of $R^{2}_{i}(t)$ for the central eleven spins within the time interval $0<t\leq 6.5$. A key observation in Fig. 4b is the pronounced dip at $\alpha=\alpha_{c}$, signifying the suppression of the impact of local quantum quench. This dip thus marks the quantum critical point characterized by enhanced quantum fluctuations, clearly delineating the boundary between the SPT phase and the trivial ordered phase. Therefore, despite the absence of local order parameters in the topological phase, our QRP, based on the local quench and local observables, demonstrates itself as a potent instrument for detecting the topological quantum phase transition.
Cluster model in a magnetic field
Finally, we extend the cluster model in Eq. (8) by introducing a transverse magnetic field:
$\mathcal{H}=\sum_{i}\left[-J_{zz}\sigma_{i}^{z}\sigma_{i+1}^{z}-h_{x}\sigma_{i}^{x}+J_{zxz}\sigma_{i}^{z}\sigma_{i+1}^{x}\sigma_{i+2}^{z}\right].$ (9)
The interplay among these three terms gives rise to a complex phase diagram [54, 55, 56, 57]: the SPT phase is stabilized when the cluster interaction term dominates with a large $J_{zxz}$, while topologically trivial phases emerge in the ferromagnetically ordered state under strong Ising interaction or in the disordered state under a strong magnetic field. Here, we fix $J_{zz}=0.1$ and vary $J_{zxz}=(1-J_{zz})\alpha$ and $h_{x}=(1-J_{zz})(1-\alpha)$, with a tuning parameter $\alpha$ adjusting the balance between the cluster and magnetic field terms. At the critical value $\alpha_{c}=0.5$, the topological quantum phase transition occurs, delineating the SPT phase ($\alpha>\alpha_{c}$) from the disordered phase ($\alpha<\alpha_{c}$[57]. Importantly, these phases are devoid of local orderings on both sides of the quantum critical point, unlike the transition examined in Fig. 4, which exhibits the local ferromagnetic order on one side. We initialize the system as
$|\Psi^{\mathrm{in}}_{k}\rangle=|{+}_{y}{+}_{y}\cdots\rangle\otimes|\psi^{\prime}(s_{k})\rangle\otimes|{+}_{y}{+}_{y}\cdots\rangle,$ (10)
where $|\psi^{\prime}(s_{k})\rangle=\sqrt{1-s_{k}}|{+}_{y}\rangle+\sqrt{s_{k}}|{-}_{y}\rangle$ and |±$|{\pm}_{y}\rangle=\left(\mid\uparrow\rangle\pm i\mid\downarrow\rangle\right)/\sqrt{2}$. $R^{2}_{i}(t)$ is subsequently evaluated from the dynamics of $\sigma_{i}^{x}$. These are chosen to clearly demonstrate the QRP by leveraging the flexibility of both the input and output.
@@
Figure 5a illustrates the estimation performance $R^{2}_{i}(t)$ under the Hamiltonian $\mathcal{H}$ for three different values of $\alpha$. In all scenarios, $R^{2}_{i}(t)$ acquires nonzero values propagating from the central site throughout the system, where the distinctions between the phases become evident. Notably, at the quantum critical point ($\alpha=\alpha_{c}$), $R^{2}_{i}(t)$ is markedly suppressed, displaying a discernibly dark spatiotemporal region in Fig. 5a. This behavior contrasts with the SPT phase at $\alpha=0.7$ and the trivial disordered phase at $\alpha=0.3$, both of which exhibits more complicated pattern with widely spread nonzero $R^{2}_{i}(t)$. To quantitatively assess these differences, we illustrate the $\alpha$ dependence of $\overline{R^{2}}$ in Fig. 5b, which is defined by averaging $R^{2}_{i}(t)$ over the central 13 sites within the time frame $0<t\leq 7$. Consistent with the previous models, $\overline{R^{2}}$ exhibits a pronounced dip at the quantum critical point, which precisely delineates the boundary between the trivial and topological phases. Indeed, the quantum fluctuations are consistently amplified near the critical point, irrespective of the system’s topological nature or the existence of local orderings. Therefore, even purely topological quantum phase transitions can be detected via the QRP, using these fluctuations in the post-local quench dynamics as witnesses.
Figure 5: Detection of the disordered-to-SPT quantum phase transition in the cluster model under a magnetic field via the QRP. a Color map of the estimation performance $R_{i}^{2}(t)$ for $\langle\sigma^{x}_{i}(t)\rangle$, in the cluster model under a magnetic field with $J_{zxz}=(1-J_{zz})\alpha$, $h_{x}=(1-J_{zz})(1-\alpha)$, and $J_{zz}=0.1$, The system size is $N=49$, and the time evolution is calculated with a bond dimension of $\chi=256$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{0<t\leq 7\}$, where $\Lambda$ contains the central 13 sites. The upper band represents the phase diagram, where the topological and disordered phases are separated by the quantum critical point at $\alpha=0.5$.
@@ DISCUSSION
To summarize, we have proposed the QRP as an versatile tool for identifying the quantum phase transitions in a variety of quantum systems, including integrable, nonintegrable, and topological systems. The QRP selectively evaluates the effects of the local quench in a manner similar to the pump-probe paradigm. Here, the parametrized local quantum quench, acting as the “pump”, injects random disturbance into the dynamics. The estimation performance of this random input parameter, through the linear transformation of the single-site operator, serves as the “probe”. This pump-probe-like process informationally discerns the underlying relationship between the local quench and the resultant dynamics of the operators. Through the QRP, we have observed that the local quench operation exerts varied influence depending on the nature of the quantum phases. At the quantum critical points, the introduced disturbance becomes subordinate to the maximally enhanced quantum fluctuations. A noticeable decline in the estimation performance observed therein serves as a marker for the transition between distinct quantum phases. These findings have established that local, nonequilibrium excitations bear the signatures of quantum phase transitions concerning global, equilibrium many-body ground states. The crux of our methodology lies in the suppression of local quenching effects on local dynamics due to amplified quantum fluctuations, the rigorous justification of which is left for future study.
@@
Our investigations have focused on 1D quantum spin systems; however, the QRP methodology is applicable to a broader range of quantum systems in more than one dimension as well. Indeed, the conformal field theories governing quantum phase transitions in two- and higher-dimensional systems exhibit significant differences from their 1D counterparts [58]. These differences manifest themselves as changes in nature of excitations, intrinsic dynamics, and critical behaviors near quantum critical points, all of which would be detectable through the QRP. Moreover, geometric frustrations in high-dimensional quantum spin systems, coupled with the inherent quantum fluctuations, give rise to a plethora of topological quantum phase transitions in a variety of systems. For instance, in certain systems, such transitions are associated with the emergence of anyons, which have promising potential for applications in topological quantum computing [59]. The examination of how anyons respond to local quantum operations through the QRP holds significant importance in light of the aforementioned applications. We leave the high-dimensional applications of the QRP for future investigations.
@@
Finally, let us remark on the design flexibility of the QRP. A fundamental advantage of employing local quantum quenches, in contrast to global ones, lies in their capability to access and elucidate quantum phenomena induced by local operations, such as specific types of excitations. When one possesses prior knowledge regarding the physics of the targeted system, especially for discerning partially understood phenomena, the optimal selections for the initial state, local operation, and read-out operators in the QRP become intuitively evident. In contrast, the true value of the QRP framework becomes markedly pronounced in the exploration of completely unknown quantum systems. Therein, the underlying physics remains undisclosed, rendering the interpretation of the complex dynamics observed a formidable task. Our QRP provides a systematic approach to this complexity, categorizing operators by meticulously investigating their dependence on various local operations through a scan of the aforementioned conditions of the QRP. Furthermore, the experimental simplicity of the QRP is noteworthy, as it requires solely single-site measurements during the early stages of the dynamics, without the need for complex quantum correlations or long-time evolutions. Consequently, we believe that the QRP presents a promising methodology for eliciting novel insights into a broad spectrum of exotic quantum many-body phenomena, extending beyond quantum phase transitions.